OpenStudy (anonymous):

What is 1/6x+5 in standard form?

OpenStudy (zzr0ck3r):

\(y=\frac{1}{6}x+5\implies y-\frac{1}{6}x=5\implies\frac{1}{6}x+(-1)y=-5\)

OpenStudy (zzr0ck3r):

@Charmy500

OpenStudy (zzr0ck3r):

can you see this?

OpenStudy (anonymous):

YES!!! I can finally see this now @zzr0ck3r thank you!!